a) Find the wavelength of the spectral line corresponding to the transition in hydrogen from then=6 state to then=3 state.
b.) Find the wavelength of the photon emitted when a hydrogen atom goes from the n=10 state to its ground state.
c) How much energy is required to remove an electron in the n-state from a hydrogen atom?
T.
1
Expert's answer
2020-05-07T10:06:50-0400

a)


"\frac{hc}{\lambda}=-13.6\left(\frac{1}{6^2}-\frac{1}{3^2}\right)"

"\frac{4.136\cdot10^{-15}}{\lambda}=-13.6\left(\frac{1}{6^2}-\frac{1}{3^2}\right)"

"\lambda=3.65\cdot10^{-15}\ m"

b)


"\frac{hc}{\lambda}=-13.6\left(\frac{1}{10^2}-\frac{1}{1^2}\right)"

"\frac{4.136\cdot10^{-15}}{\lambda}=-13.6\left(\frac{1}{10^2}-\frac{1}{1^2}\right)"

"\lambda=3.07
\cdot10^{-16}\ m"

c)


"E=\frac{13.6}{n^2}\ eV"